# Lessons Learned from a Unifying Empirical Study of Parameter-Efficient Transfer Learning (PETL) in Visual Recognition

### v2

@@ B.2.3 Selective Parameter Tuning Methods
The methods falling within this category aim to selectively update the parameters of a pre-trained model for downstream tasks. Within transfer learning, two prominent strategies, namely full fine-tuning and linear probing Kornblith et al. (2019); Zhuang et al. (2020), represent the two extremes of this category. Full fine-tuning updates all the model parameters end-to-end based on the new dataset while linear probing treats the pre-trained model as a feature extractor and only updates the prediction heads while keeping the backbone frozen. Although full fine-tuning generally exhibits superior performance compared to linear probing Zhai et al. (2019), it possesses certain limitations that may hinder its practicality in real-world production settings. Firstly, it requires running gradient descent for all parameters and necessitates storing a separate fine-tuned model for each task, incurring significant computational, memory, and storage overhead. These challenges become more salient with Transformer-based models whose parameters grow exponentially. Secondly, full fine-tuning may distort pre-trained features and underperform linear probing in out-of-distribution (OOD) scenarios Kumar et al. (2021).
To cope with the above issues, a cohort of PETL methods has emerged under this category. In addition to the two common approaches mentioned above, our investigation encompasses seven methods that can be further categorized into two groups: direct selective tuning Zaken et al. (2022); Basu et al. (2023); Xie et al. (2023), which involves the direct modification of selective weights, and efficient selective tuning Hu et al. (2021); Jie & Deng (2023); Lian et al. (2022), which approximates the weight updates with low-rank factors.
Notably, an extra advantage of methods in this category is that they introduce no additional inference latency, making them particularly favourable when inference efficiency is a priority. Methods within the direct selective tuning group abstain from introducing any new modules, thus inherently avoiding extra inference latency. Meanwhile, for methods in the efficient selective tuning group, the added modules can often be seamlessly integrated into weights of the pre-trained models through the re-parameterization techniques Jacob et al. (2018); Ding et al. (2021), thereby ensuring the absence of increased inference latency as well.
@@
Zaken et al. (2022) is a simple yet effective method that only tunes the bias parts of the pre-trained model. For each Transformer layer in ViT, BitFit updates the bias terms in the QKV projections and the FC layer in the MSA block, two FC layers in the MLP block and two LN blocks. It also updates the bias in the projection for patch embedding. The original authors underscore BitFit’s capability to achieve performance comparable to full fine-tuning or even surpass it under low and medium-data scenarios in BERT models Kenton & Toutanova (2019).
@@
Basu et al. (2023) represents another simple but strong baseline that solely tunes the two LN blocks in each Transformer layer - one before the MSA block and another before the MLP block. Given that each LN block contains merely two trainable parameters $\{\bm{W}_{LN},\bm{b}_{LN}\}\in\mathbb{R}^{D}$, LN-tune stands out as an exceedingly light-weight approach compared to other PETL methods. For instance, ViT-B/16 ($\sim$86M parameters) has only $\sim$38k LN parameters, accounting for $\sim$0.04of the total parameters.
@@
Xie et al. (2023) is a recently proposed PETL strategy designed for adapting large pre-trained diffusion models to the new domains. DiffFit exclusively fine-tunes the bias terms and the LN blocks within the network. Furthermore, it inserts learnable scale factors $\gamma$ to shift the features after the MSA and the MLP blocks, as shown in Equation 17. Consequently, DiffFit can be regarded as a combination of the BitFit and Ln-Tune, incorporating additional feature shift factors.
@@
(Low-Rank Adaptation) Hu et al. (2021) drew inspiration from recent investigations demonstrating that the learned over-parametrized models in fact reside on a low intrinsic dimension Li et al. (2018); Aghajanyan et al. (2021). Building upon this insight, the authors hypothesize that the change in weights during model adaptation also exhibits a low intrinsic rank and injects trainable low-rank decomposition matrices to approximate the weight updates. The LoRA update methodology is strategically applied to the Query/Value projection weights $\bm{W}_{Q/V}\in\mathbb{R}^{D\times D}$ within the MSA block. Concretely, the weight updates are approximated as $\bm{W}_{Q/V}+\Delta\bm{W}_{Q/V}=\bm{W}_{Q/V}+\bm{W}^{Q/V}_{\text{down}}\bm{W}^{Q/V}_{\text{up}}$ where $\bm{W}^{Q/V}_{\text{down}/\text{up}}\in\mathbb{R}^{D\times r/r\times D}$ and rank $r\ll D$. The authors use a random Gaussian initialization for $\bm{W}^{Q/V}_{\text{up}}$ and zero for $\bm{W}^{Q/V}_{\text{down}}$ so that $\Delta\bm{W}_{Q/V}=\bm{W}^{Q/V}_{\text{down}}\bm{W}^{Q/V}_{\text{up}}$ is zero at the beginning of training. The formal definition of LoRA is articulated in Equation 18, utilizing the notations delineated in Figure 9.
@@
(Factor Tuning) Jie & Deng (2023) is inspired by the recent advances in Transformer compression Wang et al. (2022a); Zhang et al. (2022a)and exploited the low-rank update paradigm (e.g., LoRA) to the extreme. While LoRA posits that the update for an individual weight matrix manifests a low-rank characteristic during fine-tuning, FacT advances the proposition that the weight updates spanning different matrices can also be effectively approximated using low-rank decomposition matrices. Specifically, FacT encapsulates the four weight matrices $\bm{W}_{Q/K/V/O}\in\mathbb{R}^{D\times D}$ in the MSA block and the two weight matrices $\bm{W}_{1}\in\mathbb{R}^{D\times 4D}$, $\bm{W}_{2}\in\mathbb{R}^{4D\times D}$ in the MLP block into a single $\bm{W}_{FacT}\in\mathbb{R}^{12M\times D\times D}$ tensor where $M$ is the number of Transformer layer. The update of $\bm{W}_{FacT}$, $\Delta\bm{W}_{FacT}$, can be decomposed into several factors to promote parameter efficiency. To this end, the authors leverage the well-established Tensor-Train (TT) Oseledets (2011)and the Tucker (TK) De Lathauwer et al. (2000) format to decompose $\Delta\bm{W}_{FacT}$. $\text{FacT}_{\text{TT}}$ and $\text{FacT}_{\text{TK}}$ are used to denote different decomposition formats for FacT and their formal definitions can be found in Equation 19.
@@
(Scale & Shift deep Features) Lian et al. (2022) employs linear transformations to adapt the intermediate features extracted by a pre-trained model. Motivated by the feature modulation methods Huang & Belongie (2017); Perez et al. (2018), SSF is designed to accommodate the distribution difference between the upstream and downstream datasets. Specifically, SSF modulates the features residing at $h_{2},h_{3},h_{5},h_{7},h_{8},h_{9}$ by incorporating scale and shift factors. To demonstrate the mechanism of SSF, let’s consider $h_{5}\in\mathbb{R}^{(N+1)\times D}$ as an illustrative example and other features can similarly undergo the same transformative process. Formally, the modulated $h_{5}$ is formulated as follows.
@@ Appendix C More Detailed Results
Learning with low-shot data is prone to over-fitting. We find that if the drop path rate — which stochastically drops a transformer block per sample Huang et al. (2016) — is set not as default (i.e., nonzero), all the methods can benefit from such a regularization. Figure 11 shows the performance gain by tuning the drop-path-rate on compared with the default 0.
Figure 11: Performance gain for PETL methods by turning drop-path-rate on.
@@
Figure 13 shows the prediction analysis discussed in section 4 for all the datasets in VTAB-1K. It is expected that their predictions are similar for datasets with very high accuracy, such as Flowers102 (avg 99.1%) and Caltech101 (avg 91.4%). Beyond them, we find that most PETL methods show diverse predictions in other datasets in VTAB-1K.
Prediction similarity within the same PETL group.
To verify if methods within the same PETL group share more prediction similarity, we plotted the prediction overlap for adapter-based methods, selective-tuning methods, and methods from different groups. As shown in Figure 12, methods within the same group share slightly more prediction similarity than those from different groups, but they still exhibit distinct predictions. Figure 4 in the main paper also supports this observation. Methods are grouped based on the categories defined in subsection 2.2. If methods within the same group had very high similarities, we would see bright squares, which are only slightly evident around BitFit, DiffFit, LayerNorm, and SSF.
@@
WiSE PETL results for all distribution shift datasets.
We provide detailed WiSE PETL performance for each distribution shift dataset in Figure 14. WiSE improves both the robustness and the in-distribution performance of PETL methods. Interestingly, even though full fine-tuning is generally less robust than PETL methods, applying WiSE allows it to achieve better performance in both target distribution and distribution shift data.
@@
Figure 14: WiSE PETL performance on all distribution shift datasets. Target distribution vs. distribution shifts
@@
This figure illustrates the relative performance compared to linear probing ($\times$) on VTAB-1K. The range between the highest and lowest accuracy across 14 PETL methods is represented by $\bullet$-$\bullet$, while ($\blacksquare$) denotes the performance of full fine-tuning.
1(b) details.
The X-axis represents the accuracy on ImageNet-1K, while the Y-axis shows the distribution shift accuracy (averaged across ImageNet-V2, ImageNet-S, ImageNet-R, and ImageNet-A). The cyan squares ($\blacksquare$) represent the zero-shot performance of the CLIP model, and stars ($\star$) denote the performance of fine-tuned models. Each curve corresponds to the WiSE+PETL method, with dots $\bullet$ indicating different mixing coefficients $\alpha$ as described in section 7.
@@
For each dataset in VTAB-1K, 15 methods (14 PETL methods plus linear probing) are ranked by accuracy. Within each dataset group (e.g.., Natural), the element $(i,j)$ in the ranking frequency matrix indicates how often method $i$ ranks $j^{th}$. For instance, in the Natural group matrix, the entry (1, 3) equals 2, meaning DiffFit ranked 3rd in two datasets within this group. The row sums correspond to the total number of datasets in each group (e.g.., 7 datasets for the Natural group). Methods are sorted by their average rank (shown in brackets), and the parameters column indicates the number of trainable parameters in millions.
Figure 4 details.
@@
Figure 4 details.
The Venn diagrams are generated by fine-tuning a pre-trained model on CIFAR100 (VTAB-1K) using LoRA, SSF, and Adapter methods. For Figure 4(a), we selected the correct predictions from the top 5K most confident samples for each method and visualized the overlap among the three methods. For Figure 4(b), we did the same for the wrong predictions, selecting from the 5K least confident samples.
@@
Different colors represent various PETL methods. Each $\bullet$ along a curve (corresponding to a single PETL method) indicates the accuracy at a specific tunable parameter size, allowing us to observe how the size of tunable parameters impacts accuracy.
@@
Each sub-figure displays accuracy on the Y-axis, with columns representing linear probing (left), the best PETL methods (middle), and full fine-tuning (right). Sub-figures (a) and (b) correspond to VTAB-1K (low-shot), while (c) and (d) correspond to many-shot settings. Different colors represent distinct datasets.

### v4

@@ B.2.3 Selective Parameter Tuning Methods
The methods falling within this category aim to selectively update the parameters of a pre-trained model for downstream tasks. Within transfer learning, two prominent strategies, namely full fine-tuning and linear probing [48, 111], represent the two extremes of this category. Full fine-tuning updates all the model parameters end-to-end based on the new dataset while linear probing treats the pre-trained model as a feature extractor and only updates the prediction heads while keeping the backbone frozen. Although full fine-tuning generally exhibits superior performance compared to linear probing [104], it possesses certain limitations that may hinder its practicality in real-world production settings. Firstly, it requires running gradient descent for all parameters and necessitates storing a separate fine-tuned model for each task, incurring significant computational, memory, and storage overhead. These challenges become more salient with Transformer-based models whose parameters grow exponentially. Secondly, full fine-tuning may distort pre-trained features and underperform linear probing in out-of-distribution (OOD) scenarios [50].
To cope with the above issues, a cohort of PEFT methods has emerged under this category. In addition to the two common approaches mentioned above, our investigation encompasses seven methods that can be further categorized into two groups: direct selective tuning [103, 7, 97], which involves the direct modification of selective weights, and efficient selective tuning [37, 44, 55], which approximates the weight updates with low-rank factors.
Notably, an extra advantage of methods in this category is that they introduce no additional inference latency, making them particularly favourable when inference efficiency is a priority. Methods within the direct selective tuning group abstain from introducing any new modules, thus inherently avoiding extra inference latency. Meanwhile, for methods in the efficient selective tuning group, the added modules can often be seamlessly integrated into weights of the pre-trained models through the re-parameterization techniques [41, 18], thereby ensuring the absence of increased inference latency as well.
@@
[103] is a simple yet effective method that only tunes the bias parts of the pre-trained model. For each Transformer layer in ViT, BitFit updates the bias terms in the QKV projections and the FC layer in the MSA block, two FC layers in the MLP block and two LN blocks. It also updates the bias in the projection for patch embedding. The original authors underscore BitFit’s capability to achieve performance comparable to full fine-tuning or even surpass it under low and medium-data scenarios in BERT models [46].
@@
[7] represents another simple but strong baseline that solely tunes the two LN blocks in each Transformer layer - one before the MSA block and another before the MLP block. Given that each LN block contains merely two trainable parameters $\{\bm{W}_{LN},\bm{b}_{LN}\}\in\mathbb{R}^{D}$, LN-tune stands out as an exceedingly light-weight approach compared to other PEFT methods. For instance, ViT-B/16 ($\sim$86M parameters) has only $\sim$38k LN parameters, accounting for $\sim$0.04of the total parameters.
@@
[97] is a recently proposed PEFT strategy designed for adapting large pre-trained diffusion models to the new domains. DiffFit exclusively fine-tunes the bias terms and the LN blocks within the network. Furthermore, it inserts learnable scale factors $\gamma$ to shift the features after the MSA and the MLP blocks, as shown in Equation 17. Consequently, DiffFit can be regarded as a combination of the BitFit and Ln-Tune, incorporating additional feature shift factors.
@@
(Low-Rank Adaptation) [37] drew inspiration from recent investigations demonstrating that the learned over-parametrized models in fact reside on a low intrinsic dimension [52, 1]. Building upon this insight, the authors hypothesize that the change in weights during model adaptation also exhibits a low intrinsic rank and injects trainable low-rank decomposition matrices to approximate the weight updates. The LoRA update methodology is strategically applied to the Query/Value projection weights $\bm{W}_{Q/V}\in\mathbb{R}^{D\times D}$ within the MSA block. Concretely, the weight updates are approximated as $\bm{W}_{Q/V}+\Delta\bm{W}_{Q/V}=\bm{W}_{Q/V}+\bm{W}^{Q/V}_{\text{down}}\bm{W}^{Q/V}_{\text{up}}$ where $\bm{W}^{Q/V}_{\text{down}/\text{up}}\in\mathbb{R}^{D\times r/r\times D}$ and rank $r\ll D$. The authors use a random Gaussian initialization for $\bm{W}^{Q/V}_{\text{up}}$ and zero for $\bm{W}^{Q/V}_{\text{down}}$ so that $\Delta\bm{W}_{Q/V}=\bm{W}^{Q/V}_{\text{down}}\bm{W}^{Q/V}_{\text{up}}$ is zero at the beginning of training. The formal definition of LoRA is articulated in Equation 18, utilizing the notations delineated in Figure 9.
@@
(Factor Tuning) [44] is inspired by the recent advances in Transformer compression [92, 106]and exploited the low-rank update paradigm (e.g., LoRA) to the extreme. While LoRA posits that the update for an individual weight matrix manifests a low-rank characteristic during fine-tuning, FacT advances the proposition that the weight updates spanning different matrices can also be effectively approximated using low-rank decomposition matrices. Specifically, FacT encapsulates the four weight matrices $\bm{W}_{Q/K/V/O}\in\mathbb{R}^{D\times D}$ in the MSA block and the two weight matrices $\bm{W}_{1}\in\mathbb{R}^{D\times 4D}$, $\bm{W}_{2}\in\mathbb{R}^{4D\times D}$ in the MLP block into a single $\bm{W}_{FacT}\in\mathbb{R}^{12M\times D\times D}$ tensor where $M$ is the number of Transformer layer. The update of $\bm{W}_{FacT}$, $\Delta\bm{W}_{FacT}$, can be decomposed into several factors to promote parameter efficiency. To this end, the authors leverage the well-established Tensor-Train (TT) [72]and the Tucker (TK) [14] format to decompose $\Delta\bm{W}_{FacT}$. $\text{FacT}_{\text{TT}}$ and $\text{FacT}_{\text{TK}}$ are used to denote different decomposition formats for FacT and their formal definitions can be found in Equation 19.
@@
(Scale & Shift deep Features) [55] employs linear transformations to adapt the intermediate features extracted by a pre-trained model. Motivated by the feature modulation methods [39, 73], SSF is designed to accommodate the distribution difference between the upstream and downstream datasets. Specifically, SSF modulates the features residing at $h_{2},h_{3},h_{5},h_{7},h_{8},h_{9}$ by incorporating scale and shift factors. To demonstrate the mechanism of SSF, let’s consider $h_{5}\in\mathbb{R}^{(N+1)\times D}$ as an illustrative example and other features can similarly undergo the same transformative process. Formally, the modulated $h_{5}$ is formulated as follows.
@@ Appendix C More Detailed Results
Learning with low-shot data is prone to over-fitting. We find that if the drop path rate — which stochastically drops a transformer block per sample [38] — is set not as default (i.e., nonzero), all the methods can benefit from such a regularization. Figure 11 shows the performance gain by tuning the drop-path-rate on compared with the default 0.
Figure 11: Performance gain for PEFT methods by turning drop-path-rate on.
@@
Figure 13 shows the prediction analysis discussed in section 4 for all the datasets in VTAB-1K. It is expected that their predictions are similar for datasets with very high accuracy, such as Flowers102 (avg 99.1%) and Caltech101 (avg 91.4%). Beyond them, we find that most PEFT methods show diverse predictions in other datasets in VTAB-1K.
Prediction similarity within the same PEFT group.
To verify if methods within the same PEFT group share more prediction similarity, we plotted the prediction overlap for adapter-based methods, selective-tuning methods, and methods from different groups. As shown in Figure 12, methods within the same group share slightly more prediction similarity than those from different groups, but they still exhibit distinct predictions. 3(a) in the main paper also supports this observation. Methods are grouped based on the categories defined in subsection 2.2. If methods within the same group had very high similarities, we would see bright squares, which are only slightly evident around BitFit, DiffFit, LayerNorm, and SSF.
@@
WiSE PEFT results for all distribution shift datasets.
We provide detailed WiSE PEFT performance for each distribution shift dataset in Figure 14. WiSE improves both the robustness and the in-distribution performance of PEFT methods. Interestingly, even though full fine-tuning is generally less robust than PEFT methods, applying WiSE allows it to achieve better performance in both target distribution and distribution shift data.
Performance comparison between DINOv2 and IN21k.
To compare the performance of DINOv2 and IN21k, we selected several PEFT methods and datasets from VTAB-1K. The results presented in Table 7 reveal several interesting findings:
(1) Improved Linear Probing Performance: Linear probing generally shows improved results with DINOv2, indicating that its extracted features are more robust and discriminative than those from IN21k.
(2) Deteriorated Full Fine-Tuning Performance: Conversely, full fine-tuning performance significantly worsens with DINOv2, suggesting that models fully fine-tuned on DINOv2 are more susceptible to overfitting.
(3) Adapter-Based Methods Performance: Among the three adapter-based methods evaluated—Houl. Adapter, AdaptFormer, and Convpass—we observe performance enhancements in most datasets for AdaptFormer and Convpass. In contrast, Houl. Adapter exhibits significant degradation across all datasets. This disparity may be attributed to architectural differences: AdaptFormer and Convpass insert their adapter modules in parallel with existing modules such as Multi-Head Self-Attention (MSA) and/or Multi-Layer Perceptron (MLP), whereas Houl. Adapter inserts its adapter sequentially after the MSA and MLP layers. We hypothesize that the sequential design of Houl. Adapter leads to more substantial alterations of intermediate features compared to the parallel design, potentially explaining the observed decrease in performance.
Linear Full SSF Houl. Adapter Adapt- Former Convpass LoRA
Natural Caltech101 Dinov2 89.8 83.2 90.3 22.1 92.5 91.7 92.3
IN21k 86.6 89.9 89.8 92.1 91.8 92.1 92.6
$\Delta$ 3.2 -6.7 0.5 -70.0 0.7 -0.4 -0.3
DTD Dinov2 74.9 45.2 77.0 14.6 78.8 77.0 78.4
IN21k 65.7 61.9 68.8 72.3 70.5 72.0 69.8
$\Delta$ 9.2 -16.7 8.2 -57.7 8.3 5.0 8.6
Pets Dinov2 93.4 68.7 92.6 7.1 94.0 92.3 94.1
IN21k 89.3 85.8 91.4 91.7 91.8 91.3 90.5
$\Delta$ 4.1 -17.1 1.2 -84.6 2.2 1.0 3.6
Sun397 Dinov2 55.1 23.9 52.5 2.6 56.5 56.1 55.7
IN21k 53.2 36.8 56.5 55.4 56.7 55.9 55.5
$\Delta$ 1.9 -12.9 -4.0 -52.8 -0.2 0.2 0.2
Specialized Camelyon Dinov2 83.2 77.9 83.9 77.1 84.4 86.3 85.4
IN21k 83.1 81.6 86.1 88.7 86.8 87.7 87.5
$\Delta$ 0.1 -3.7 -2.2 -11.6 -2.4 -1.4 -2.1
EuroSAT Dinov2 89.2 66.9 93.9 60.2 93.8 93.8 94.2
IN21k 90.0 88.1 94.5 95.3 95.0 95.8 94.9
$\Delta$0.8 -21.2 -0.6 -35.1 -1.2 -2.0 -0.7
Resisc45 Dinov2 78.8 25.9 82.0 24.5 88.6 87.6 84.2
IN21k 74.9 81.6 83.2 86.5 86.5 85.9 85.9
$\Delta$ 3.9 -55.7 -1.2 -62.0 2.1 1.7 -1.7
Retinopathy Dinov2 75.3 73.6 76.0 73.6 76.0 76.0 75.5
IN21k 74.6 73.6 74.8 75.2 76.3 75.9 75.7
$\Delta$ 0.7 0.0 1.2 -1.6 -0.3 0.1 -0.2
Structured Clevr-Count Dinov2 47.5 27.3 71.2 38.1 91.2 87.8 89.8
IN21k 37.5 56.2 80.1 82.9 82.9 82.3 82.9
$\Delta$ 10.0 -28.9 -8.9 -44.8 8.3 5.5 6.9
DMLab Dinov2 44.1 30.7 51.8 39.1 51.8 53.1 54.5
IN21k 36.5 48.2 53.0 53.8 52.8 53.8 51.8
$\Delta$ 7.6 -17.5 -1.2 -14.7 -1.0 -0.7 2.7
KITTI Dinov2 60.3 47.1 81.0 46.8 83.4 82.6 83.8
IN21k 64.6 77.9 81.4 79.6 80.0 78.1 79.9
$\Delta$4.3 -30.8 -0.4 -32.8 3.4 4.5 3.9
dSpr-Ori Dinov2 47.2 17.5 56.1 10.0 57.9 55.6 57.2
IN21k 29.4 46.6 52.1 54.3 53.0 55.3 47.2
$\Delta$ 17.8 -29.1 4.0 -44.3 4.9 0.3 10.0
Table 7: Performance comparison between DINOv2 and IN21k.
@@
Figure 14: WiSE PEFT performance on all distribution shift datasets. Target distribution vs. distribution shifts
@@
This figure illustrates the relative performance compared to linear probing ($\times$) on VTAB-1K. The range between the highest and lowest accuracy across 14 PEFT methods is represented by $\bullet$-$\bullet$, while ($\blacksquare$) denotes the performance of full fine-tuning.
1(c) details.
The X-axis represents the accuracy on ImageNet-1K, while the Y-axis shows the distribution shift accuracy (averaged across ImageNet-V2, ImageNet-S, ImageNet-R, and ImageNet-A). The cyan squares ($\blacksquare$) represent the zero-shot performance of the CLIP model, and stars ($\star$) denote the performance of fine-tuned models. Each curve corresponds to the WiSE+PEFT method, with dots $\bullet$ indicating different mixing coefficients $\alpha$ as described in section 7.
@@
For each dataset in VTAB-1K, 15 methods (14 PEFT methods plus linear probing) are ranked by accuracy. Within each dataset group (e.g., Natural), the element $(i,j)$ in the ranking frequency matrix indicates how often method $i$ ranks $j^{th}$. For instance, in the Natural group matrix, the entry (1, 3) equals 2, meaning DiffFit ranked 3rd in two datasets within this group. The row sums correspond to the total number of datasets in each group (e.g., 7 datasets for the Natural group). Methods are sorted by their average rank (shown in brackets), and the parameters column indicates the number of trainable parameters in millions.
3(a) details.
@@
3(b) details.
The Venn diagrams are generated by fine-tuning a pre-trained model on CIFAR100 (VTAB-1K) using LoRA, SSF, and Adapter methods. For 3(b)(a), we selected the correct predictions from the top 5K most confident samples for each method and visualized the overlap among the three methods. For 3(b)(b), we did the same for the wrong predictions, selecting from the 5K least confident samples.
@@
Different colors represent various PEFT methods. Each $\bullet$ along a curve (corresponding to a single PEFT method) indicates the accuracy at a specific tunable parameter size, allowing us to observe how the size of tunable parameters impacts accuracy.
@@
Each sub-figure displays accuracy on the Y-axis, with columns representing linear probing (left), the best PEFT methods (middle), and full fine-tuning (right). Sub-figures (a) and (b) correspond to VTAB-1K (low-shot), while (c) and (d) correspond to many-shot settings. Different colors represent distinct datasets.